# Long-range Entanglement and Role of Realistic Interaction in Braiding of Non-Abelian Quasiholes in Fractional Quantum Hall Phases

### v2

@@ Acknowledgement
The braiding phase can be extracted from the Berry phase data. In general, when a quasihole moves in a closed circle on the sphere, there exists three contributions to the Berry phase: the Aharonov-Bohm phase, the parallel transport phase (coming from the curvature of the sphere), and the braiding phase (coming from any other quasihole presence within the closed loop). The first two terms are both proportional to the solid angle $\Omega$ enclosed by the loop and only the braiding phase is $\Omega$-independent. Thus, we can extract the braiding phase by determining the $\Omega$-independent part of the phase by linear fitting. There is however a caveat: since quasiholes are objects with a finite size, this is only true if they are sufficiently far apart from each other. Thus, in practice, in a plot of the Berry phase against $\Omega$, we find that only the portion of the graph at larger $\Omega$ can be fitted with a line (see Fig. 6 in main text). However, in most cases with larger systems, the braiding phase can be reliably extracted from the $y$-intercept of the linear fit of the right portion of the graph. In Fig. S4, one also finds that the qualitative features of the graphs are quite consistent across system sizes
Figure S4: Left column: Berry phase as a function of solid angle for a potential consisting solely of three pinning potentials (a) 8 electrons (c) 9 electrons (e) 10 electrons (f) 11 electrons. right column: Berry phase as a function of solid angle for a Hamiltonian consisting of a pinning potential and an addition $\hat{V}_{1}^{2bdy}$ (b) 8 electrons (d) 9 electrons (f) 10 electrons (h) 11 electrons. In each figure, the exchange phase (in multiple of $2\pi$) is given by the $y$-intercept in the fitting equation indicated on the plot title.
S6 Tuning the non-Abelian gap in presence of impurity
@@
where $\hat{c}^{\dagger}_{m}$ and $\hat{c}_{m}$ respectively creates and annihilates an electron at orbital $m$, and $\eta_{m_{1}m_{2}}$ are complex numbers whose real and imaginary parts are each a random number between $-1$ and $1$, with a constraint that $\eta_{m_{2}m_{1}}=\eta_{m_{1}m_{2}}^{*}$. The overall strength of this random potential is modulated by a real scalar factor $v_{i}$ in Eq. (S39) below. In general, this overall factor cannot be too large, otherwise the random potential will dominate over any interaction in the system[69]. The total toy-model Hamiltonian used in this section is

### v3

@@ Acknowledgement
The braiding phase can be extracted from the Berry phase data. In general, when a quasihole moves in a closed circle on the sphere, there exists three contributions to the Berry phase: the Aharonov-Bohm phase, the parallel transport phase (coming from the curvature of the sphere), and the braiding phase (coming from any other quasihole presence within the closed loop). The first two terms are both proportional to the solid angle $\Omega$ enclosed by the loop and only the braiding phase is $\Omega$-independent. Thus, we can extract the braiding phase by determining the $\Omega$-independent part of the phase by linear fitting. There is however a caveat: since quasiholes are objects with a finite size, this is only true if they are sufficiently far apart from each other. Thus, in practice, in a plot of the Berry phase against $\Omega$, we find that only the portion of the graph at larger $\Omega$ can be fitted with a line (see Fig. 7 in the main text). However, in most cases with larger systems, the braiding phase can be reliably extracted from the $y$-intercept of the linear fit of the right portion of the graph.
S11 Tuning the non-Abelian gap in realistic systems
S11.1 Balancing $\hat{V}_{1}^{2bdy}$ and $\hat{V}_{3}^{2bdy}$
Figure S3: The non-Abelian gap as a function of interpolation between $\hat{V}_{1}^{2bdy}$ ($\lambda=0$) to $\hat{V}_{3}^{2bdy}$ ($\lambda=1$) (see Eqs. (S73) and (S74)) for (a) $N_{e}=6$ electrons and (b) $N_{e}=8$ electrons.
As discussed in the main text, the non-Abelian gap refers to the amount of energy required for a pair of $\psi$-anyons to spontaneously transmute to a pair of $1$-anyons. Ensuring that this energy is zero (or as low as possible) is crucial in observing non-Abelian statistics. In numerics, we study this gap by working on a MR system with four quasiholes in the even sector and apply a two-body interaction as well as two pinning potentials:
$\mathcal{H}=\lambda_{0}\hat{V}^{3bdy}+\hat{V}^{2bdy}+v_{0}\hat{V}_{\text{pins}}$ (S73)
Setting $\lambda_{0}\to\infty$ ensures the system stays in the MR CHS, while two potential pins specified by $\hat{V}_{\text{pins}}$ ensures only two quasiholes (each with charge $e/2$) are present in the system. If the two pins are far enough from each other (one at each pole on the sphere), we expect the ground state to be either two $1$-type quasiholes or two $\psi$-type quasiholes, depending on the exact form of the two-body interaction $\hat{V}^{2bdy}$. Since $\hat{V}_{1}^{2bdy}$ and $\hat{V}_{3}^{2bdy}$ each prefers a different type of quasiholes (namely, the $\psi$-type and $1$-type, respectively), it is enough to consider a toy model that interpolates between the two:
$\hat{V}^{2bdy}=(1-\lambda)\hat{V}_{1}^{2bdy}+\lambda\hat{V}_{3}^{2bdy}$ (S74)
In general, one expect the gap between the two lowest energy to close as $\lambda$ is tuned from 0 to 1, as the ground state transmutes from two $\psi$-anyons into two $1$-anyons. This is what we see in Fig 8 of the main text for $N_{e}=10$ electrons. Here we present more numerical results to show that this feature is robust across many system sizes (see Fig. S3). While the exact value of $\lambda$ at which the gap attains a minimum value changes, it is true across all system sizes that there always exists an interaction interpolating between $\hat{V}_{1}^{2bdy}$ and $\hat{V}_{3}^{2bdy}$ where the non-Abelian gap is minimized.
S11.2 In presence of impurities
@@
where $\hat{c}^{\dagger}_{m}$ and $\hat{c}_{m}$ respectively creates and annihilates an electron at orbital $m$, and $\eta_{m_{1}m_{2}}$ are complex numbers whose real and imaginary parts are each a random number between 1-1 and $1$, with a constraint that $\eta_{m_{2}m_{1}}=\eta_{m_{1}m_{2}}^{*}$. The overall strength of this random potential is modulated by a real scalar factor $v_{i}$ in Eq. (S76) below. In general, this overall factor cannot be too large, otherwise the random potential will dominate over any interaction in the system. The total toy-model Hamiltonian used in this section is
@@
S11.3 Sample thickness and screening as tuning parameters
Using the same principle, we can also investigate the effect of sample thickness and screening in tuning the non-Abelian gap. As suggested earlier, increasing the thickness increases the effective interaction range, thereby effectively increasing the $\hat{V}_{3}^{2bdy}$-to-$\hat{V}_{1}^{2bdy}$ ratio. On the other hand, increasing the screening strength reduces the effective interaction range, thereby effectively reducing this ratio. However, any realistic interaction contains not only $\hat{V}_{1}^{2bdy}$ and $\hat{V}_{3}^{2bdy}$, but also higher terms $\hat{V}_{5}^{2bdy}$, $\hat{V}_{7}^{2bdy}$, and so on. These terms may have unpredictable effect on the anyons dynamics.
Here we study the effect of these parameter in tuning the non-Abelian gap by the following toy Hamiltonian
$\mathcal{H}=\lambda_{0}\hat{V}^{3bdy}+\hat{V}^{2bdy}+v_{0}\hat{V}_{\text{pins}}+v_{i}\hat{V}_{\text{rand}}^{1bdy}$ (S77)
where $\hat{V}^{2bdy}$ is either the ZDS interaction $\hat{V}_{\text{ZDS}}^{\text{(n)}}$ or the Yukawa potential $\hat{V}_{\text{Yukawa}}^{\text{(n)}}$. Here, the superscript (n) denotes that these interactions are normalized, i.e. they are scaled by a factor such that all pseudopotential coefficients sum up to 1. This is to ensure that any decrease in the gap is not due to the decrease in overall strength of the interaction.
The results for preliminary calculations with $N_{e}=6$ electrons are shown in Fig. S5 which shows the non-Abelian gap decreasing with layer thickness, as expected from increasing the interaction range. On the other hand, the non-Abelian gap does not vary monotonously with screening strength. It does not close exactly, but there is a screening strength at which it achieves a minimum. While these results show a proof of concept that it is possible to tune the non-Abelian gap with experimental parameter, to realize this idea requires further extensive numerics with experimental input. An important thing to note that in order to observe non-Abelian statistics, any interaction that minimizes non-Abelian gap must not also destroy the FQH plateau (this is not a concern in our numerics because we have set $\lambda_{0}\to\infty$ in Eq. (S77)).
Figure S5: (a) The non-Abelian gap as a function of layer thickness with two-body interaction modelled by the ZDS interaction (b) The non-Abelian gap as a function of screen strength with the screened Coulomb interaction modelled by the Yukawa potential. Both are calculated for systems with $N_{e}=6$ electrons.